OpenStudy (anonymous):

what are the steps to solve for x in this problem: 2x - y= 3/4x + 6

OpenStudy (compassionate):

\[2x -y = \frac{3}{4x} + 6\] See that y? Yea, put it on the right side by adding See that x? Get it on the left side by multiplying.

OpenStudy (anonymous):

then what

OpenStudy (compassionate):

Well - for starters, can you show me what you have so far?

OpenStudy (anonymous):

3/4 x 2x = y+6

OpenStudy (compassionate):

Hey, I am actually very tired. You can plug this into a calculator and get the answer?

OpenStudy (anonymous):

yeah thx for your help

OpenStudy (compassionate):

Use this: http://www.mathway.com/problem.aspx?p=algebra